OpenStudy (anonymous):

Henry’s law constant for carbon dioxide at 38°C is 0.00228 mol L-1 atm-1. Calculate the pH of a solution of carbon dioxide at 38°C in equilibrium with the gas at a partial pressure of 3.20 atm.

OpenStudy (anonymous):

First you need to use, well, Henry's law to find the solubility of the gas in liquid: C = k * P\(_{gas}\) Where: C = solubility of a gas in a particular solvent at a fixed temperature (in \(\frac{mol}{L}\), or simply M) k = 0.00228 (in \(\frac{mol}{L\cdot atm}\)) P\(_{gas}\) = partial pressure of the gas (in atm) Next you need to consider this equilibrium reaction: CO(\_2\) (g) + H\(2\)O (l) \( \leftarrow \rightarrow \) H\(_2\)CO\(_3\) (aq) and this: pH = -log[H\(_3\)0\(^+\)] where the number of ions is proportional to the acid presen and then finally you need the experimental values for pKw @ 38\(^o\)C and pKa for carbonic acid (called the acid dissociation constant) http://research.chem.psu.edu/brpgroup/pKa_compilation.pdf http://en.wikipedia.org/wiki/Acid_dissociation_constant You should now have every thing you need to answer this question :-)

OpenStudy (anonymous):

Thank you!